OpenStudy (anonymous):

Manny and Emma are both competing in a long jump competition at their school. Manny can jump four fifths of the distance that Emma can jump. Manny can jump 28 inches. How far can Emma jump?

OpenStudy (anonymous):

\[28=\frac{ 4 }{ 5 }x\] Emma is x. Solve

OpenStudy (anonymous):

m = manny jumping e = emma jumping then m = (4/5) * e m = 28 then 28 = (4/5) * e solve for emma, so divide both sides by (4/5) 28/(4/5) = e using fractions in division is like this: (28/1) / (4/5) is the same as (28/1) * (5/4) = (28*5)/4 so e = (28*5)/4